For the reaction 2A --> B + ½ C, answer the following questions:

a.If –rA = 10 mol L-1 min-1, calculate the rate of reaction with respect to Band C, rB and rC, respectively.

b.10 moles of A are fed to a reactor. If 70% conversion occurs, calculate the moles of all chemical species. Are the total moles conserved during the course of reaction?

c.The reaction order with respect to A is 2.0. Provide the units that would be possessed by the rate constant, in terms of moles, liters, and minutes.
Answer #1

a) The reaction rates are calculated:

rB = 10 M A / min * (1 M B / 2 M A) = 5 M B / min

rC = 10 M A / min * (1/2 M C / 2 M A) = 2.5 M C / min

b) The moles of each substance are calculated:

n A = (1-0.7) * 10 = 3 mol

n B = (10 - 3) mol A * (1 mol B / 2 mol A) = 3.5 mol B

n C = (10 - 3) mol A * (1/2 mol C / 2 mol A) = 1.75 mol C

c) The units will be:

[k] = L mol-1 min-1

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.